OpenStudy (goformit100):

Twice the sum of a number and 4 is equal to the sum of triple the number and 8.

OpenStudy (anonymous):

2(x + 4) = 3x + 8

OpenStudy (ganpat):

2(x+4) = 3x + 8..

OpenStudy (jiteshmeghwal9):

let the no. be 'x' . then 2(x+4)=3x+8

OpenStudy (jiteshmeghwal9):

solve for 'x' :)

OpenStudy (anonymous):

Do you need to find the number? Then just solve 2(x+4) = 3x + 8 distribute the 2 combine like terms solve for x

OpenStudy (goformit100):

Thanks

OpenStudy (jiteshmeghwal9):

this question is wrong is becoz 2(x+4)=3x+8 \(2x+8 \ne\ 3x+8 \)